OpenStudy (anonymous):

help? square roots are so confusing. 3 sqrt 7 (1 - sqrt 7)

OpenStudy (anonymous):

\[3\sqrt{7} (1-\sqrt{7})\]

OpenStudy (solomonzelman):

\[3\sqrt{7}\times1+3\sqrt{7}\times (-\sqrt{7})\]

OpenStudy (solomonzelman):

\[3\sqrt{7}\times (-\sqrt{7}) --->3(-1)(\sqrt{7)}(\sqrt{7)}=?\]

OpenStudy (solomonzelman):

\[-3\sqrt{7}^2 --->(-3\times7)--->(-21)\]

OpenStudy (solomonzelman):

\[3\sqrt{7}-21\]

OpenStudy (anonymous):

do i finish that?

OpenStudy (solomonzelman):

this is in the most simplified form